OpenStudy (anonymous):

Which of the following equations is of an ellipse with x-intercepts of (3, 0) and (-3, 0) and y-intercepts at (0, 1) and (0, -1)?

OpenStudy (apoorvk):

Hmm, so, looks like: Now for such an ellipse with x-intercepts at (a,0) and (-a,0) and y-intercepts at (0,b) and (0,-b), the equation representing the ellipse is:\[\large \frac{x^2}{a^2}+\frac{y^2}{b^2} = 1\] Use this, can you?

OpenStudy (amistre64):

intercepts are the same no matter what the form of the equation. when x=0, you get the y intercept when y=0, you get the x intercepts

OpenStudy (apoorvk):

aye^

OpenStudy (anonymous):

thanks yall